### v2

@@ Methods
We employ an A-inspired 41 global greedy tree search approach to explore synthesis pathways for a target molecule. This method leverages the template prioritizer models to identify the most applicable reaction rules, which are then applied to the target molecule to derive precursor molecules.
In this search framework, precursor molecules are represented as nodes in a tree, while reactions and their associated conditions serve as edges. However, not all precursors may be readily available or affordable. We define “buyable” as being available for purchase at a cost of under \$100/g.
The search algorithm proceeds as follows:
@@
Algorithm 1 outlines the global greedy tree search algorithm.
@@
2:$N_{0}\leftarrow$ initial chemical structure with cost $C_{0}$
3:$PQ\leftarrow\{\}$triangleright Priority queue
4:add $N_{0}$ to $PQ$ with score $P_{0}=\infty$
@@
6:     $N\leftarrow$ pop node with highest score from $PQ$
7:     if goal_check($N$then
@@
9:     $R\leftarrow$ find_rules($N$)
@@
11:         $N^{\prime}\leftarrow$ apply_rule($r$, $N$)
@@
17:         add $N^{\prime}$ to subtrees of $N$
18:         insert $N^{\prime}$ into $PQ$ with priority $P^{\prime}$
19:function find_rules($N$)
20:     return predicted rules for $N$
@@
22:     return new chemical structure from applying $r$ to $N$
@@
24:     $P^{\prime}\leftarrow f(C^{\prime},T^{\prime},S^{\prime})$triangleright Compute score
@@
26:function goal_check($N$)
27:     return True if $N$ meets goal conditions (low $C$ or max depth), else False
@@
Scoring
We introduce a user-tunable scoring methodology to evaluate synthesis pathways.
The first feature of the score is the cost. We utilize the ASKCOS buyable dataset 42, Molport 43, Mcule 44, and Chemspace 45 APIs with the ChemPrice library 46 to determine the cost and availability of molecules. A molecule is considered buyable if it can be purchased for less than \$100/g. Users have the flexibility to adjust this threshold based on their objectives. The score is calculated as $-\frac{\text{price}}{500}$, as we aim to maximize the score, thus exploring reactions involving cheaper precursors first. The value 500 is chosen as a normalizer because we define anything costing above \$500/g as non-buyable; the user can modify this threshold. Our collaborating experimental chemists verified these initial thresholds.
The second feature of the score is the reaction temperature. We employ a CASP tool developed by Gao et al. 47 to predict the temperature at which a reaction might occur. This prediction is essential for exploring pathways where reactions may not have been previously documented or experimented on. To find the predicted temperature, we take the weighted average of the top ten predictions made by the tool, a strategy suggested by the authors to improve accuracy. The score is calculated as $-\frac{\text{temperature}}{300}$, as we aim to maximize the score, thus exploring reactions involving lower temperatures first. The value 300 is chosen as a normalizer because we define any reaction requiring over 300 C as non-practical; the user can modify this threshold.
The third feature of the score is the solvent and reagent greenness-toxicity score. We use the same CASP tool developed by Gao et al. 47 to predict the solvent and reagent required for a reaction. We curate a toxicity and greenness dataset, assigning a score of -1 for toxic molecules, 0 for neutral molecules, and +1 for green or natural molecules. We use the ACS Solvent Selection Guide 48 to classify 100 commonly used solvents. We also use the SuperNatural 3.0 dataset 49, containing 350,000 natural products, and the T3DB dataset 50 to classify 4,000 toxic molecules.
These individual scores are summed based on user-tunable weights, allowing users to prioritize features according to their preferences during the pathway search.
Results
We use the enzymatic reaction dataset by curating the BKMS dataset 33 and the Rhea database 34, and the synthetic reaction dataset from the USPTO repository 36 to train and test our template prioritizers.
Model Performance
Accuracy of literature rule Avg. number of applicable rules Presence of applicable rule
@@ Results
DNN 5 0.100 0.312 0.442 0.485 0.098 0.898 4.876 14.005 0.098 0.539 0.900 0.973
MHN 23 0.181 0.566 0.760 0.809 0.099 1.004 5.647 16.388 0.099 0.531 0.861 0.967
Ours (Enz) 0.183 0.576 0.763 0.814 0.101 1.076 6.125 17.380 0.101 0.553 0.884 0.976
@@
Table 1: Performance Metrics for Different Models on Template Prioritization. The table is divided into three sets of metrics. The first set (columns 2-5) shows the accuracy of the presence of the literature rule in the top predictions (T1, T10, T50, T100). The second set (columns 6-9) represents the average number of applicable rules in the top predictions. The third set (columns 10-13) indicates the accuracy of the presence of at least one applicable rule in the top predictions. $\dagger$ The synthetic model is trained and tested on a separate, larger dataset, while the other three models are trained and tested on the same dataset.
@@
    Accuracy of the presence of the literature rule in the top n predictions (T1, T10, T50, T100): This metric evaluates the model’s ability to prioritize reaction templates documented in the literature. The “literature rule” refers to the reaction template extracted from the corresponding reaction in our dataset. This metric is crucial for assessing how well the model replicates known synthetic pathways. While this metric is important for evaluating the replication of known synthetic pathways, assessing the model’s ability to propose novel and feasible pathways is equally critical. Therefore, we also consider the following metrics, which provide a more comprehensive view of the model’s performance. This is the commonly used metric in recent works 5, 23.
@@
    Average number of applicable rules in the top n predictions: This metric provides insight into the diversity and feasibility of the proposed reaction templates. A rule is considered “applicable” when the RDKit library successfully applies the transformation to the target molecule, resulting in valid precursor structures. This metric is particularly important as it indicates the model’s capability to explore broader chemical space and suggest alternative synthetic routes that may not be present in the literature but are chemically plausible.
@@
    Accuracy of the presence of at least one applicable rule in the top n predictions: This metric complements the other two by measuring the model’s practical utility in retrosynthetic analysis. It assesses how often the model suggests at least one viable synthetic step, ensuring the progression of retrosynthetic analysis in chemically plausible directions.
@@
Our enzymatic model (Ours (Enz)) demonstrates slightly better performance across most metrics compared to the baseline DNN (Deep Neural Network) 5 and MHN 23 models. Notably, for the accuracy of presence of literature rule T1 metric, our model achieves an accuracy of 18.3%, surpassing the best performing DNN (10%) and slightly outperforming the state-of-the-art MHN (18.1%). This improvement is consistent across all the metrics, indicating improved template prioritization capabilities with higher number and accuracy of applicable rules on average.
The synthetic ensemble model exhibits exceptional performance, with a T1 accuracy of 42.2%, significantly outperforming all other models. This ensemble approach proves particularly effective in exploring diverse synthetic pathways, as evidenced by the high average number of applicable rules in the top predictions. For instance, when applying the top 10 rules from this model, an average of 6.315 rules are applicable, providing a rich set of precursors for further exploration by applying just 10 rules. This high applicability rate is crucial for generating diverse and feasible synthetic routes, enabling chemists to consider multiple pathways and optimize their synthesis strategies based on various factors such as cost, yield, or environmental impact.
To manage the computational demands of processing millions of reactions while preserving the diversity of the chemical space, we partitioned the synthetic reactions into five equally sized subsets via a scaffold split. This approach reduces memory overhead and speeds up training by ensuring that each subset is sufficiently small to handle yet still representative of the entire dataset. Notably, a single synthetic template prioritizer trained on one subset achieves a top-1 accuracy of 36.6%, whereas the ensemble of all five models yields an improved top-1 accuracy of 42.2%.
Comparison with Pathways from the Literature
@@
We assessed our synthesis planning framework by benchmarking it against established datasets and frameworks, including PaRoutes 31 and ChemByDesign 32. Precisely, we assess the pathway lengths generated by our framework and report metrics on the number of replicated pathways and the average number of predicted pathways per compound. We also present some case studies involving a hybrid approach combining enzymatic and synthetic pathways in Figure 2(a) and Supplementary Figures 12 and 14. This comparative analysis aimed to assess our framework’s accuracy, efficiency, and versatility in predicting viable synthetic routes for complex target molecules.
@@
PaRoutes 31 is a robust benchmarking framework for evaluating multi-step retrosynthesis methods. It comprises two datasets of 10,000 synthetic routes derived from the patent literature alongside a curated list of purchasable molecules and reactions suitable for training retrosynthesis models. For this study, we utilized a scaffold split technique to select pathways associated with the ten most commonly occurring scaffolds in the PaRoutes dataset, which are given in Supplementary Table 3. This approach enabled us to efficiently assess our framework’s ability to predict synthetic routes that align with those documented in the patent literature while ensuring chemical diversity.
@@
ChemByDesign 32, on the other hand, is an online platform that organizes experimentally verified reaction pathways by name, year, and author. To ensure an unbiased evaluation, we focused on pathways discovered after 2021 that are under six steps long. These pathways were intentionally excluded from our training dataset to test the predictive capabilities of our framework on novel and unexplored synthetic routes. This strategy provided a unique opportunity to evaluate how effectively our framework generalizes to unseen data.
@@
The results of this comparative analysis are summarized in Figure 2. Figure 2(a) presents a representative tree of reaction pathways generated by our framework for a target molecule in the PaRoutes dataset. Notably, our predicted pathway utilizes inexpensive precursors and environment-friendly solvents, such as ethanol and methanol, demonstrating its potential for sustainable synthesis planning. The tree also highlights alternative routes that leverage naturally occurring molecules while avoiding toxic solvents like dichloromethane (DCM). These results underscore the framework’s ability to prioritize green chemistry principles without compromising synthetic feasibility and cost.
Figure 2(b) provides quantitative performance metrics for our framework compared to PaRoutes and ChemByDesign. The first bar plot illustrates the number of molecules attempted and successfully solved by each method. Our framework solved twelve out of fifteen molecules attempted. Furthermore, our framework replicated nine known pathways from the literature while suggesting a high number of pathways on average per molecule (4.8 and 2.2). These results demonstrate its ability to explore diverse chemical spaces and propose multiple viable options for synthetic planning.
Finally, the third bar plot compares the lengths of predicted pathways relative to known literature routes. Our framework identified thirty six shorter or equivalent-length pathways, with two cases [^1] N-(3-[5-Chloro-2-(difluoromethoxy)phenyl]-1-2-[(4-pyridinylmethyl)amino]ethyl-1H-pyrazol-4-yl)pyrazolo[1,5-a]pyrimidine-3-carboxamide and 4-(1-tert-butyl-4-oxo-5H-pyrazolo[4,3-c]pyridin-3-yl)thiophene-2-carboxamide yielding routes two steps shorter than those documented in PaRoutes and one case (Lupinine) yielding a route two steps shorter than that documented in ChemByDesign. This capability to optimize pathway length is particularly valuable in industrial settings where shorter synthetic routes can lead to significant cost savings and improved process efficiency.
Supplementary Figure 9 illustrates an alternative pathway for Lupinine 51 (found in ChemByDesign) synthesis predicted by MHNpath. Our framework identified a streamlined three-step route with moderate reaction conditions ($12.58\degree$C, $11.1\degree$C, and $-23.81\degree$C) and low precursor costs (\$87.36/g and \$0.10/g), contrasting with the five-step approach by Wang et al. that spans temperatures from $-78\degree$C to $85\degree$C.
@@
Figure 2: (a) Tree of reaction pathways. The tree shows a representative example for a pathway presented in PaRoutes 31. Our predicted pathway is producible using cheap precursors and less toxic, naturally occurring solvents.
@@
Figure 3: (a-g) Published pathways. Overview of existing pathways to produce dronabinol. (h) Previously predicted pathways by Levin et al. 5. Four-step reaction pathway to produce dronabinol as predicted by Levin et al. 5. (i) Our predicted pathway. Three-step reaction to produce dronabinol from cheap precursors in ambient temperatures. We also replicated some of the other pathways. (j) Performance metrics for comparison with other models. These plots present the number of molecules solved, the average number of pathways predicted, and the distribution of predicted pathway lengths 5, 27.
Comparison with Other Models

### v3

@@ 2.3   Searching algorithm
We employ an A-inspired Hart1968 global greedy tree search approach to explore synthesis pathways for a target molecule. This method leverages the template prioritizer models to identify the most applicable reaction rules, which are then applied to the target molecule to derive precursor molecules. In this search framework, precursor molecules are represented as nodes in a tree, while reactions and their associated conditions serve as edges. However, not all precursors may be readily available or affordable. We define "buyable” as being available for purchase at a cost of under \$100/g. The search algorithm proceeds as follows:
@@
Algorithm˜1 outlines the global greedy tree search algorithm.
@@
[…]
4:add $N_{0}$ to PQPQ with score $P_{0}=\infty$
@@
6:  $N\leftarrow$ pop node with highest score from PQPQ
7:  if goal_check(NN) then
@@
9:  $R\leftarrow$ find_rules(NN)
@@
11:   $N^{\prime}\leftarrow$ apply_rule($r$, NN)
@@
17:   add $N^{\prime}$ to subtrees of NN
18:   insert $N^{\prime}$ into PQPQ with priority $P^{\prime}$
19:function find_rules(NN)
20:  return predicted rules for NN
@@
22:  return new chemical structure from applying $r$ to NN
@@
24:  $P^{\prime}\leftarrow f(C^{\prime},T^{\prime},S^{\prime})$ $\triangleright$ Compute score
@@
26:function goal_check(NN)
27:  return True if NN meets goal conditions (low $C$ or max depth), else False
@@
2.4   Scoring system
We introduce a user-tunable scoring methodology to evaluate synthesis pathways. In this approach individual scores for three features are obtained and summed based on user-tunable weights, allowing users to prioritize features according to their preferences during the pathway search. The first feature of the score is the cost. We utilize the ASKCOScoley2019robotic buyable dataset, Molport MolPort2020, Mcule Kiss2012, and Chemspace chemspace APIs with the ChemPriceSorkun2024 library to determine the cost and availability of molecules. A molecule is considered buyable if it can be purchased for less than \$100/g levin2022merging, sterling2015zinc. Users have the flexibility to adjust this threshold based on their objectives. The score is calculated as $\frac{\text{price}}{500}$, as we aim to maximize the score, thus exploring reactions involving cheaper precursors first. The value 500 is chosen as a normalizer because we define anything costing above \$500/g as non-buyable; the user can modify this threshold. These threshold aligns with the affordability criteria for research-grade building blocks utilized in major chemical databases sterling2015zinc.
The second feature of the score is the reaction temperature. We employ a CASP tool developed by Gao et al. gao2018using to predict the temperature at which a reaction might occur. This prediction is essential for exploring pathways where reactions may not have been previously documented or experimented on. To find the predicted temperature, we take the weighted average of the top ten predictions made by the tool, a strategy suggested by the authors to improve accuracy. The score is calculated as $\frac{\text{temperature}}{300}$, as we aim to maximize the score, thus exploring reactions involving lower temperatures first. The value 300 is chosen as a normalizer because we define any reaction requiring over 300 C as non-practical; this cutoff reflects the upper operating limits of standard synthetic laboratory equipment razzaq2010continuous.
The third feature of the score is the solvent and reagent greenness-toxicity score. We use the same CASP tool developed by Gao et al. gao2018using to predict the solvent and reagent required for a reaction. We curate a toxicity and greenness dataset, assigning a score of -1 for toxic molecules, 0 for neutral molecules, and +1 for green or natural molecules. We use the ACS Solvent Selection Guide ACSGreenChemistryInstitute2011 to classify 100 commonly used solvents. We also use the SuperNatural 3.0 dataset 10.1093/nar/gkac1008, containing 350,000 natural products, and the T3DB dataset 10.1093/nar/gku1004 to classify 4,000 toxic molecules.
3    Results and Discussion
In this section, we evaluate the performance of the MHNpath framework in prioritizing reaction templates and generating feasible retrosynthetic pathways. We benchmark our predictive models against established baselines and analyze the practical utility of the proposed scoring system in guiding pathway discovery.
3.1   Model performance
Lit. Rule Accuracy Avg. Applicable Rules Presence of Applicable Rule
@@ 3.1   Model performance
DNN levin2022merging 0.100 0.312 0.442 0.485 0.128 0.898 4.876 14.005 0.128 0.539 0.900 0.973
MHNreact seidl2022improving 0.181 0.566 0.760 0.809 0.199 1.004 5.647 16.388 0.199 0.531 0.861 0.967
Ours (Enz) 0.183 0.576 0.763 0.814 0.201 1.076 6.125 17.380 0.201 0.553 0.884 0.976
@@
Table 1: Performance Metrics for Template Prioritization. The best, second best, and third best results for the enzymatic dataset are color coded. The table is divided into three sets of metrics. The first set (columns 2-5) shows the accuracy of the presence of the literature rule in the top predictions (T1, T10, T50, T100). The second set (columns 6-9) represents the average number of applicable rules in the top predictions. The third set (columns 10-13) indicates the accuracy of the presence of at least one applicable rule in the top predictions.
$\dagger$ The synthetic model is trained and tested on a separate, larger dataset, while the other three models are trained and tested on the same dataset.
@@
    Accuracy of the presence of the literature rule in the top n predictions (T1, T10, T50, T100): This metric evaluates the model’s ability to prioritize reaction templates documented in the literature. The "literature rule" refers to the reaction template extracted from the corresponding reaction in our dataset. This metric is crucial for assessing how well the model replicates known synthetic pathways. This is the commonly used metric in recent works levin2022merging, seidl2022improving. While this metric is important for evaluating the replication of known synthetic pathways, assessing the model’s ability to propose novel and feasible pathways is equally critical. Therefore, we also consider the following metrics, which provide a more comprehensive view of the model’s performance.
@@
    Average number of applicable rules in the top n predictions: This metric provides insight into the diversity and feasibility of the proposed reaction templates. A rule is considered "applicable" when the RDKit library successfully applies the transformation to the target molecule, resulting in valid precursor structures. This metric is particularly important as it indicates the model’s capability to explore broader chemical space and suggest alternative synthetic routes that may not be present in the literature but are chemically plausible. This is critical for our tree search, as a higher density of valid rules in the top predictions (T50, T100) expands the branching factor of the search tree with high-quality and newer candidates.
@@
    Accuracy of the presence of at least one applicable rule in the top n predictions: This metric complements the other two by measuring the model’s practical utility in retrosynthetic analysis. It assesses how often the model suggests at least one viable synthetic step, ensuring the progression of retrosynthetic analysis in chemically plausible directions. High scores here ensure the search algorithm rarely hits "dead ends".
@@
Our enzymatic model (Ours (Enz)) demonstrates good performance across most metrics compared to the baseline DNN (Deep Neural Network) levin2022merging and MHN seidl2022improving models. It is important to contextualize the baseline performance, the DNN achieves a Top-1 accuracy of only 10%, which underscores the inherent complexity of the retrosynthetic prediction task. Unlike standard classification problems, template prioritization involves selecting the correct chemical transformation from a massive search space of valid reaction templates. In this high-dimensional context, low absolute accuracy scores are standard, and incremental gains represent significant practical improvements.
First, compared to the standard DNN baseline levin2022merging, our model delivers a large improvement. The DNN achieves a Top-1 literature accuracy of only 10% and a Top-1 applicability presence of 12.8%. Our model nearly doubles these figures to 18.3% and 20.1%, respectively. Given the high-dimensional search space (selecting 1 out of >>17,000 templates), this jump represents a significant improvement in the model’s ability to prioritize relevant chemistry. This trend is mirrored in the applicability metrics, indicating that our model not only retrieves ground truth better but also generates a higher number of valid chemical precursors.
Second, while our architecture shares similarities with the MHN model seidl2022improving, our specific enhancements (Xavier initialization, rigorous dropout, and hyperparameter tuning) yield better convergence and robustness, particularly when predicting precursors for structurally complex or challenging target molecules. For example, in the Avg. Number of Applicable Rules metric at T50, our model outperforms the MHN baseline (6.125 vs. 5.647). This indicates that our model retrieves approximately 8.5% more valid chemical options in the top-50 candidates. This increased density of applicable rules provides the global greedy search algorithm with a richer pool of precursors, reducing the likelihood of missing a viable pathway.
Finally, the synthetic ensemble model (Ours (Syn)) demonstrates exceptional performance on the USPTO dataset, achieving a T1 accuracy of 42.2%. Most notably, it exhibits a massive applicability rate: the Top-10 predictions yield, on average, 6.315 valid reaction rules. This high applicability ensures that the synthetic branch of our hybrid planner can almost always identify multiple feasible precursors, allowing the scoring system to aggressively optimize for cost and environmental impact without running out of chemical options.
It is important to note that the baseline DNN levin2022merging, the MHN model seidl2022improving, and our enzymatic model (Ours (Enz)) were trained and evaluated on the exact same enzymatic dataset splits, ensuring a direct and fair comparison.
3.2   Comparison with pathways from the literature
@@
We assessed our synthesis planning framework by benchmarking it against established datasets and frameworks, including PaRoutes D2DD00015F and ChemByDesign draghici2012chemistry. Precisely, we assess the pathway lengths generated by our framework and report metrics on the number of replicated pathways and the average number of predicted pathways per molecule. We also present some case studies involving a hybrid approach combining enzymatic and synthetic pathways in Figure˜2(a) and Supplementary Figures 12 and 14. This comparative analysis aimed to assess our framework’s accuracy, efficiency, and versatility in predicting viable synthetic routes for complex target molecules.
@@
PaRoutes D2DD00015F is a robust benchmarking framework for evaluating multi-step retrosynthesis methods. It comprises two datasets of 10,000 synthetic routes derived from the patent literature alongside a curated list of purchasable molecules and reactions suitable for training retrosynthesis models. For this study, we utilized a scaffold split technique to select pathways associated with the ten most commonly occurring scaffolds in the PaRoutes dataset, which are given in Supplementary Table 3. This approach enabled us to efficiently assess our framework’s ability to predict synthetic routes that align with those documented in the patent literature while ensuring chemical diversity.
@@
ChemByDesign draghici2012chemistry, on the other hand, is an online platform that organizes experimentally verified reaction pathways by name, year, and author. To ensure an unbiased evaluation, we focused on pathways discovered after 2021 that are under six steps long. These pathways were intentionally excluded from our training dataset to test the predictive capabilities of our framework on novel and unexplored synthetic routes. This strategy provided a unique opportunity to evaluate how effectively our framework generalizes to unseen data.
@@
The results of this comparative analysis are summarized in Figure 2. Figure˜2(a) presents a representative tree of reaction pathways generated by our framework for a target molecule in the PaRoutes dataset. This evaluation was conducted on an expanded test set comprising 130 molecules from the PaRoutes dataset and 5 novel targets from ChemByDesign. Notably, our predicted pathway utilizes inexpensive precursors and environment-friendly solvents, such as ethanol and methanol, demonstrating its potential for sustainable synthesis planning. The tree also highlights alternative routes that leverage naturally occurring molecules while avoiding toxic solvents like dichloromethane (DCM); further examples of these green alternatives are illustrated in Supplementary Figure 9. These results underscore the framework’s ability to prioritize green chemistry principles without compromising synthetic feasibility and cost.
@@
Figure 2: (a) Tree of reaction pathways. The tree shows a representative example for a pathway presented in PaRoutes D2DD00015F. Our predicted pathway is producible using cheap precursors and less toxic, naturally occurring solvents.
@@
Figure˜2(b) provides quantitative performance metrics for our framework compared to PaRoutes and ChemByDesign. The first bar plot illustrates the number of molecules attempted and successfully solved by each method. Our framework solved twelve out of fifteen molecules attempted. Furthermore, our framework replicated nine known pathways from the literature while suggesting a high number of pathways on average per molecule (4.8 and 2.2). These results demonstrate its ability to explore diverse chemical spaces and propose multiple viable options for synthetic planning.
Finally, the third bar plot compares the lengths of predicted pathways relative to known literature routes. Our framework identified thirty six shorter or equivalent-length pathways, with two casesΩ yielding routes two steps shorter than those documented in PaRoutes and one case (Lupinine) yielding a route two steps shorter than that documented in ChemByDesign. This capability to optimize pathway length is particularly valuable in industrial settings where shorter synthetic routes can lead to significant cost savings and improved process efficiency. $\Omega$: N-(3-[5-Chloro-2-(difluoromethoxy)phenyl]-1-2-[(4-pyridinylmethyl)amino]ethyl-1H-pyrazol-4-yl)pyrazolo[1,5-a]pyrimidine-3-carboxamide and 4-(1-tert-butyl-4-oxo-5H-pyrazolo[4,3-c]pyridin-3-yl)thiophene-2-carboxamide}
Supplementary Figure 9 illustrates an alternative pathway for Lupinine wang2024enantioselective (found in ChemByDesign) synthesis predicted by MHNpath. Our framework identified a streamlined three-step route with moderate reaction conditions ($2.58\degreeC, 11$$1.1\degreeC$, and $23.81\degreeC$) and low precursor costs (\$87.36/g and \$0.10/g), contrasting with the five-step approach by Wang et al. that spans temperatures from $78\degreeC$ to $85\degreeC$. The test set of utilized from PaRoutes has been provided in the Supplementary Data.
3.3   Comparison with other models